# Keep the Cost Down: A Review on Methods to Optimize LLM’s KV-Cache Consumption

### v2

@@ Abstract
Large Language Models (LLMs), epitomized by ChatGPT’s release in late 2022, have revolutionized various industries with their advanced language comprehension. However, their efficiency is challenged by the Transformer architecture’s struggle with handling long texts. KV-Cache has emerged as a pivotal solution to this issue, converting the time complexity of token generation from quadratic to linear, albeit with increased GPU memory overhead proportional to conversation length. With the development of the LLM community and academia, various KV-Cache compression methods have been proposed. In this review, we dissect the various properties of KV-Cache and elaborate on various methods currently used to optimize the KV-Cache space usage of LLMs. These methods span the pre-training phase, deployment phase, and inference phase, and we summarize the commonalities and differences among these methods. Additionally, we list some metrics for evaluating the long-text capabilities of large language models, from both efficiency and capability perspectives. Our review thus sheds light on the evolving landscape of LLM optimization, offering insights into future advancements in this dynamic field. Links to the papers mentioned in this review can be found in our Github Repo https://github.com/zcli-charlie/Awesome-KV-Cache.
@@ 1 Introduction
Since the release of ChatGPT, Large Language Models (LLMs) are gradually having a profound impact on people’s lives and are standing out in various fields (Wu et al., 2023; Roumeliotis & Tselikas, 2023b; a). These LLMs face a computational challenge: their Decoder-Only Transformer architecture has a quadratic time complexity when understanding text sequences. During inference, the auto-regressive decoding mechanism amplifies this issue, as it repeats the process for each token generated. KV-Cache, by storing the keys and values tensor in attention module generated by past tokens, can reduce the time complexity required to generate each token to linear, greatly improving inference efficiency. However, the use of KV-Cache is not all advantageous. KV-Cache will increase linearly with the length of the sequence, and the memory required will become larger and larger, especially for giant models like GPT-3 (Floridi & Chiriatti, 2020). Moreover, each individual dialogue needs to save its own KV-Cache, and different dialogues can hardly reuse them, which will become a bottleneck in the generation speed on modern inference hardware like GPU as they usually suffers from the low memory bandwidth comparing to their computing speed (Yu et al., 2022).
@@
Recent months have seen emerging work on optimizing KV-Cache, making it a critical focus for enhancing LLMs’ performance with longer contexts. This review presents various methods of KV-Cache optimization, clarifying their interrelationships and comparing their core ideas. We discuss extensive methods to reduce memory space: before inference, the model itself can be compressed, or the architecture can be changed, completely abandoning attention with quadratic complexity; during inference, from the model input prompt level, to the embedding level, and then to the KV-Cache level, compression can be performed. This paper primarily focuses on editing, modifying, and optimizing KV-Cache itself, although other methods are briefly mentioned in Appendix C. These optimizations are considered the safest, most effective, and compatible approach known to date. A general preview can be found in Figure 1.
@@
This review unfolds in chronological order of the large model, from the training phase, to the deployment phase, and finally to the post-training phase. In the training stage, we will introduce the KV-Cache compression methods that can be used during model pre-training. These methods are usually the most effective, but they are not suitable for modifying existing models or scenarios with low computational power. In the deployment stage, we will introduce the use of different frameworks to optimize the use of KV-Cache. The methods in this section will not make a large number of modifications to KV-Cache itself, but can significantly optimize its efficiency in the same environment. Finally, in the post-training stage, we will introduce a large number of on-time optimization methods for KV-Cache, mainly including two methods: Eviction and Quantization.
@@
Additionally, we introduce metrics to assess LLMs’ performance on long texts, vital for KV-Cache optimization. These metrics are divided into efficiency and performance aspects. Efficiency measures include model generation speed and space occupancy improvement, while performance metrics evaluate the impact on model capabilities, ensuring any capability loss remains acceptable.
@@
In conclusion, this review chronologically introduces KV-Cache optimization methods in LLMs, aiming to enhance model inference efficiency and context length. It seeks to advance LLMs to serve humans more effectively, efficiently, and sustainably.
Figure 1: An overview of the main structure of KV-Cache compression methods
@@ 2 Preliminary and Notations
Before introducing the KV-Cache compression method, we must rigorously define frequently occurring symbols to avoid ambiguity in the narrative. The core function of LLM is to transform a natural number sequence input into equal amount of multiple probability distributions $\mathcal{LLM}(X)=\mathcal{LLM}(x_{1}x_{2}\dots x_{n})=p_{1}p_{2}\dots p_{n}$. In this paper, we denote $X$ as the input sequence of integers, named tokens, with the length $n$, composed by $x_{i}$ as the $i$-th token. Then LLM translates them into $p_{i}$, which is the probability of the next token corresponding to $x_{i}$. In a LLM, the entire set of possible values for $x_{i}$ constitutes the vocabulary, $\mathcal{V}$. The size of the vocabulary is denoted as $V=|\mathcal{V}|$. Naturally, we have $X\in\mathcal{V}^{n}$, $x_{i}\in\mathcal{V}=\{0,1,\dots,V-1\}$, and $p_{i}\in\mathbb{R}^{V}$. The input $X$ is translated from the natural language input PROMPT in tokenization, and usually $p_{n}$ is used to predict the next token with various sampling algorithms when auto-regression decoding.
@@
Each Transformer Decoder Block consists of two parts: Self-Attention and Feed-Forward Network (FFN). Each part has its own residual connection and Layer Normalization (Norm) operation. In this paper, we focus only on the Self-Attention part. In the Decoder Block, each $h_{i}^{(l)}$ is mapped to three new vectors $q_{i}^{(l)}$, $k_{i}^{(l)}$, and $v_{i}^{(l)}$ using the trainable matrices $W_{q}^{(l)}$, $W_{k}^{(l)}$, paired with position embedding matrix $R_{i}$, and $W_{v}^{(l)}$. In other words, $H^{(l)}$ is mapped to $Q^{(l)}$, $K^{(l)}$, and $V^{(l)}$. Another matrix $W_{o}^{(l)}$ is then used to map the output of the Self-Attention layer to the input ${h^{\prime}}_{i}^{(l)}$ for a single token, and ${H^{\prime}}^{(l)}$ for the entire sequence. The process is given by Formula 1.
${H^{\prime}}^{(l)}=W_{o}^{(l)}\cdot\mathrm{softmax}\left(\mathrm{mask}\left(\frac{Q^{(l)}\cdot{K^{(l)}}^{\dagger}}{\sqrt{d}}\right)\right)\cdot V^{(l)}$ (1)
Modern LLMs utilize multi-head attention (MHA). The idea is to split $q$, $k$, and $v$ into $n_{h}$ smaller blocks, named heads, each with $d_{h}=d/n_{h}$ dimensions. For the $j$-th head of $l$-th layer, we denote $\{Q,K,V,W_{q},W_{k},W_{v},q_{i},k_{i},v_{i}\}^{(l,j)}$ for the corresponding partition of these matrices or vectors. In MHA, Causal Mask mask()($\cdot$)( ⋅ ) was used to prevent earlier tokens attend on later ones, by adding $-\mathrm{inf}$ to the upper right triangle of the pre-softmax attention matrix. This crucial property ensures that the $K$ and $V$ computed by preceding tokens will not be affected by subsequent tokens. Therefore, for each token newly added during auto-regression decoding, we only need to update previous $K$ and $V$, as $\{K^{(l)},k_{n}^{(l)}\}\longrightarrow K^{(l)}$ and $\{V^{(l)},v_{n}^{(l)}\}\longrightarrow V^{(l)}$, and perform Formula 2. The $K$ and $V$ which we retained are called KV-Cache.
${h_{n}^{\prime}}^{(l)}=W_{o}^{(l)}\cdot\mathrm{Concatenate}_{j=1}^{n_{h}}\left(\mathrm{softmax}\left(\mathrm{mask}\left(\frac{q_{n}^{(l,j)}\cdot{K^{(l,j)}}^{\dagger}}{\sqrt{d_{h}}}\right)\right)\cdot V^{(l,j)}\right)$ (2)
@@ 3 Training Stage Optimization
For LLMs that adopt the traditional Decoder-Only Transformer architecture, the most effective KV-Cache compression method emerges during the pre-training phase. This is because, in this phase, the model possesses the greatest plasticity. The primary adjustment in this phase is to the model architecture, which, while still retaining the excellent properties of Attention, reduces the size of the generated Keys and Values vectors to a quarter or even less.
Shazeer (2019) proposed Multi-Query Attention (MQA) based on Multi-Head Attention (MHA). Shazeer claimed that even if we retain only one head for keys and values, we can still achieve good model performance. In this case, different query heads calculate attention scores with the same key head. Although there is only one head left for value, it will receive $n_{h}$ different combination weights, resulting in $n_{h}$ combinations. Doing so can instantly optimize the KV-Cache space usage to $1/n_{h}$ of the original.
Reducing the number of keys and values heads from $n_{h}$ to $1$ is undoubtedly an aggressive strategy. While it optimizes performance, it inevitably sacrifices some performance. Hence, Ainslie et al. (2023) proposed GQA, a method that can better balance speed and performance. In GQA, all query heads no longer calculate attention scores with the same key head, but are divided into $n_{g}$ groups. Each group calculates attention scores with one key head, serving as the $n_{h}/n_{g}$ different combination weights for the corresponding value head in this group. All $n_{h}$ results are then concatenated. Ultimately, the KV-Cache we need to store will be reduced to $n_{g}/n_{h}$. Compared to MHA and MQA, GQA introduces an adjustable parameter $n_{g}$. When $n_{g}=1$, we get MQA, which is aggressively efficiency-oriented; when $n_{g}=n_{h}$, we get MHA, which is aggressively performance-oriented. When $n_{g}$ is between $1$ and $n_{h}$, efficiency and performance achieve a more granular balance. Among the popular open-source models, LLaMA2-70B (Touvron et al., 2023) uses GQA with $n_{g}=2$, Mistral (Jiang et al., 2023a) with $n_{g}=8$, and Yi (AI et al., 2024) with $n_{g}=4$. The boldest model is DeciLM-7B (Team, 2023), which uses different $n_{g}$ values on different layers: the first few layers are $4$, then decay to $2$, then to $1$, and finally return to $4$ on the last layer. A comparison of these three methods can be found at Figure 2.
In addition to saving KV-Cache space, using MQA and GQA has another advantage: it can save a substantial number of parameters within the Attention module, with the ratio of savings $\eta=0.5+0.5\cdot(1+n_{g}/n_{h})$. With an equal amount of parameters, models using these techniques will have more parameters left for the feed-forward layer. The current usage of GQA and MQA in open-source models can be found in Appendix  A.
@@
Yen et al. (2024) proposed a new idea for the LLM architecture aside from GQA. The authors proposed CEPE, a framework that combines the pre-trained LLM with an Encoder module that serves as context compressor. In the RAG scenario, the ultra-long context will greatly increase the need to store KV-Cache. CEPE’s additional Encoder compresses these reference texts, and then inputs them into the Decoder through cross-attention. The compressed text has a shorter sequence length than before, so the KV-Cache is saved. This method can use the pre-trained LLM, but the added decoder and cross-attention layer still require a lot of extra training.
@@ 4 Deploy-Stage Optimization
In the use of KV-Cache to accelerate LLM inference, with the generation of each new token, the operations of $\{K,k_{n}\}\longrightarrow K$ and $\{V,v_{n}\}\longrightarrow V$ are repeatedly executed. This means that the space occupied by Keys and Values in memory will continue to increase, but each time it only increases a little, which is a horrible property. For an un-optimized inference system, a large number of repeated operations to create and release space in memory will be performed, and the same data will be meaninglessly copied multiple times. This process is time-consuming and may cause the production of a massive amount of memory fragments that are difficult to recycle. Therefore, an excellent inference system, specifically designed for the high-frequency and multiple small growth properties of KV-Cache, is an important way to improve the efficiency of KV-Cache.
@@
Kwon et al. (2023) introduced the Paged Attention mechanism and the vLLM framework. Paged Attention draws on the page memory mechanism widely used in CPU memory, and uses an additional mapping table to map the KV-Cache that used to be stored continuously to discontinuous GPU memory. During inference, in order to generate the next token, it is necessary to call KV-Cache for calculation, and this process can also be efficiently completed through a set of custom CUDA kernels. The use of the Paged Attention mechanism results in almost no unused memory fragments and efficient inference.
@@
Lin et al. (2024) further developed the idea of Paged Attention. Through the DistAttention it proposed and the DistKV-LLM built on it, KV-Cache was able to achieve distributed deployment on multiple servers. This significantly improved the efficiency of providing LLM services using large-scale cloud servers.
@@
Ye et al. (2024) applied a property of KV-Cache to reuse KV-Cache between different dialogues, achieving acceleration of the pre-fill stage and optimization of GPU memory occupancy. By establishing a dictionary tree for all historical dialogues, and finding the longest common prefix in the dictionary tree when a new dialogue request is received and reusing the corresponding KV-Cache for this part, the ChunkAttention made the model avoid repeated calculation of some tokens in the pre-fill stage, speeding up the response speed of the deployment system. When providing services for larger-scale requests, especially when there is a long consistent system prompt across every request, this optimization method can significantly reduce the time required from the start of calculating the request to returning the first token. In addition, this method can also be quickly combined with Paged Attention mentioned above, because all that needs to be done is to create a few pointers, and Paged Attention can just adapt to this.
Jin et al. (2024) proposed InfLLM, a method that allows large models to achieve near-infinite context without additional training and uses very little additional KV-Cache. This idea is based on the Sliding Window Attention mechanism (SWA). For KV-Cache outside the window, it is not completely discarded, but is relegated to CPU memory. All Values are relegated, while some representative parts of Keys are retained. During generation, the new query calculates attention with the Keys inside the window and the earlier retained Keys. Only when an earlier Key obtains a higher attention score will the Keys and Values blocks in the CPU memory be activated.

### v3

# Keep the Cost Down: A Review on Methods to Optimize LLM’s KV Cache Consumption

@@ Abstract
Large Language Models (LLMs), epitomized by ChatGPT’s release in late 2022, have revolutionized various industries with their advanced language comprehension. However, their efficiency is challenged by the Transformer architecture’s struggle with handling long texts. KV Cache has emerged as a pivotal solution to this issue, converting the time complexity of token generation from quadratic to linear, albeit with increased GPU memory overhead proportional to conversation length. With the development of the LLM community and academia, various KV Cache compression methods have been proposed. In this review, we dissect the various properties of KV Cache and elaborate on various methods currently used to optimize the KV Cache space usage of LLMs. These methods span the pre-training phase, deployment phase, and inference phase, and we summarize the commonalities and differences among these methods. Additionally, we list some metrics for evaluating the long-text capabilities of large language models, from both efficiency and capability perspectives. Our review thus sheds light on the evolving landscape of LLM optimization, offering insights into future advancements in this dynamic field. Links to the papers mentioned in this review can be found in our Github Repo https://github.com/zcli-charlie/Awesome-KVCache.
@@ 1 Introduction
Since the release of ChatGPT, Large Language Models (LLMs) are gradually having a profound impact on people’s lives and are standing out in various fields (Wu et al., 2023; Roumeliotis & Tselikas, 2023b; a). These LLMs face a computational challenge: their Decoder-Only Transformer architecture has a quadratic time complexity when processing text sequences. During inference, the auto-regressive decoding mechanism amplifies this issue, as it repeats the process for each token generated. KV Cache, by storing the keys and values tensor in attention module generated by past tokens, can reduce the time complexity required to generate each token to linear, greatly improving inference efficiency. However, the use of KV Cache is not all advantageous. KV Cache will increase linearly with the length of the sequence, and the memory required will become larger and larger, especially for giant models like GPT-3 (Floridi & Chiriatti, 2020). Moreover, each individual dialogue needs to save its own KV Cache, and different dialogues can hardly reuse them, which will become a bottleneck in the generation speed on modern inference hardware like GPU as they usually suffers from the low memory bandwidth comparing to their computing speed (Yu et al., 2022).
@@
Recent months have seen emerging work on optimizing KV Cache, making it a critical focus for enhancing LLMs’ performance with longer contexts. This review presents various methods of KV Cache optimization, clarifying their interrelationships and comparing their core ideas. We discuss extensive methods to reduce memory space: before inference, the model itself can be compressed, or the architecture can be changed, completely abandoning attention with quadratic complexity; during inference, from the model input prompt level, to the embedding level, and then to the KV Cache level, compression can be performed. This paper primarily focuses on editing, modifying, and optimizing KV-Cache itself, with other methods briefly mentioned in Appendix C. These optimizations are considered the safest, most effective, and compatible approach known to date. A general preview can be found in Figure 1.
@@
This review unfolds in chronological order of the LLMs, from the training phase, to the deployment phase, and finally to the post-training phase. In the training stage, we will introduce the KV Cache compression methods that can be used during model pre-training. These methods are usually the most effective, but they are not suitable for modifying existing models or scenarios with low computational budget. In the deployment stage, we will introduce the use of different frameworks to optimize the use of KV Cache. The methods in this section will not make a large number of modifications to KV Cache itself, but can significantly optimize its efficiency in the same environment. Finally, in the post-training stage, we will introduce a large number of on-time optimization methods for KV Cache, mainly including three methods: Eviction, Merging, and Quantization.
@@
Additionally, we introduce metrics to assess LLMs’ performance on long texts, which are vital for KV Cache optimization. These metrics are divided into efficiency and performance aspects. Efficiency measures include model generation speed and space occupancy improvement, while performance metrics evaluate the impact on model capabilities, ensuring any capability loss remains acceptable.
@@
In conclusion, this review chronologically introduces KV Cache optimization methods in LLMs, aiming to enhance model inference efficiency and context length. This review aims to summarize and analyze various LLM optimization methods from the perspective of KV-Cache, enabling the readers to better understand KV Cache optimization and build more effective, efficient, and sustainable LLMs.
Figure 1: An overview of the main structure of KV Cache compression methods
@@ 2 Preliminary and Notations
Before introducing the KV Cache compression method, we must rigorously define frequently occurring symbols to avoid ambiguity in the narrative. The core function of LLM is to transform a natural number sequence input into equal amount of multiple probability distributions $\mathcal{LLM}(X)=\mathcal{LLM}(x_{1}x_{2}\dots x_{n})=p_{1}p_{2}\dots p_{n}$. In this paper, we denote $X$ as the input sequence of integers, named tokens, with the length $n$, composed by $x_{i}$ as the $i$-th token. Then LLM translates them into $p_{i}$, which is the probability of the next token corresponding to $x_{i}$. In a LLM, the entire set of possible values for $x_{i}$ constitutes the vocabulary, $\mathcal{V}$. The size of the vocabulary is denoted as $V=|\mathcal{V}|$. Naturally, we have $X\in\mathcal{V}^{n}$, $x_{i}\in\mathcal{V}=\{0,1,\dots,V-1\}$, and $p_{i}\in\mathbb{R}^{V}$. The input $X$ is translated from the natural language input PROMPT in tokenization, and usually $p_{n}$ is used to predict the next token with various sampling algorithms when auto-regression decoding.
@@
Each Transformer Decoder Block consists of two parts: Self-Attention and Feed-Forward Network (FFN). Each part has its own residual connection and Normalization (Norm) operation. In this paper, we only focus on the Self-Attention part. In the Decoder Block, each $h_{i}^{(l)}$ is mapped to three new vectors $q_{i}^{(l)}$, $k_{i}^{(l)}$, and $v_{i}^{(l)}$ using three trainable matrices $W_{q}^{(l)}$, $W_{k}^{(l)}$, paired with position embedding matrix $R_{i}$, and $W_{v}^{(l)}$. In other words, $H^{(l)}$ is mapped to $Q^{(l)}$, $K^{(l)}$, and $V^{(l)}$. Another matrix $W_{o}^{(l)}$ is then used to map the output of the Self-Attention layer to the input ${h^{\prime}}_{i}^{(l)}$ for a single token, and ${H^{\prime}}^{(l)}$ for the entire sequence. The process is given by Formula 1.
${H^{\prime}}^{(l)}=\mathrm{softmax}\left(\mathrm{mask}\left(\frac{Q^{(l)}\cdot{K^{(l)^{\mathrm{T}}}}}{\sqrt{d}}\right)\right)\cdot V^{(l)}\cdot W_{o}^{(l)}$ (1)
Modern LLMs utilize multi-head attention (MHA). The idea is to split $q$, $k$, and $v$ into $n_{h}$ smaller blocks, named heads, each with $d_{h}=d/n_{h}$ dimensions. For the $j$-th head of $l$-th layer, we denote $\{Q,K,V,W_{q},W_{k},W_{v},q_{i},k_{i},v_{i}\}^{(l,j)}$ for the corresponding partition of these matrices or vectors. In MHA, Causal Mask mask()($\cdot$)( ⋅ ) was used to prevent earlier tokens attend on later ones, by adding $-\mathrm{inf}$ to the upper right triangle of the pre-softmax attention matrix. This crucial property ensures that the $K$ and $V$ computed by preceding tokens will not be affected by subsequent tokens. Therefore, for each token newly added during auto-regression decoding, we only need to update previous $K$ and $V$, as $\{K^{(l)},k_{n}^{(l)}\}\longrightarrow K^{(l)}$ and $\{V^{(l)},v_{n}^{(l)}\}\longrightarrow V^{(l)}$, and perform Formula 2. The $K$ and $V$ which we retained are called KV Cache.
${h_{n}^{\prime}}^{(l)}=\mathrm{Concatenate}_{j=1}^{n_{h}}\left(\mathrm{softmax}\left(\mathrm{mask}\left(\frac{q_{n}^{(l,j)}\cdot{K^{(l,j)^{\mathrm{T}}}}}{\sqrt{d_{h}}}\right)\right)\cdot V^{(l,j)}\right)\cdot W_{o}^{(l)}$ (2)
@@ 3 Training Stage Optimization
For LLMs that adopt the traditional Decoder-Only Transformer architecture, the most effective KV Cache compression method emerges during the pre-training phase. This is because, in this phase, the model possesses the greatest plasticity. The primary adjustment in this phase is to the model architecture, which, while still retaining the excellent properties of Attention, reduces the size of the generated Keys and Values vectors to a quarter or even less.
Shazeer (2019) proposed Multi-Query Attention (MQA) based on Multi-Head Attention (MHA). Shazeer claimed that even if we retain only one head for keys and values, we can still achieve satisfactory model performance. In this case, different query heads calculate attention scores with the same key head. Although there is only one head left for value, it will receive $n_{h}$ different combination weights, resulting in $n_{h}$ combinations. Doing so can instantly optimize the KV Cache space usage to $1/n_{h}$ of the original.
Reducing the number of keys and values heads from $n_{h}$ to $1$ is undoubtedly an aggressive strategy. Hence, Ainslie et al. (2023) proposed Grouped Query Attention (GQA), a method that can better balance speed and performance. In GQA, query heads are divided into $n_{g}$ groups. Each group shares one key head, serving as the $n_{h}/n_{g}$ different combination weights for the corresponding value head. Ultimately, the KV Cache we need to store will be reduced to $n_{g}/n_{h}$. Compared to MHA and MQA, GQA introduces an adjustable parameter $n_{g}$. When $n_{g}=1$, we get MQA; when $n_{g}=n_{h}$, we get MHA. When $n_{g}$ is between $1$ and $n_{h}$, efficiency and performance achieve a more granular balance. Moreover, GQA can save a substantial number of parameters within the Attention module, with the ratio of savings $\eta=0.5+0.5\cdot n_{g}/n_{h}$. A comparison of MHA, MQA and GQA can be found at Figure 2. The current usage of GQA and MQA in open-source models can be found in Appendix  A.
@@
If GQA reuses the KV Cache within layers, then the method proposed by Sun et al. (2024); Brandon et al. (2024); Goldstein et al. (2024) involves inter-layer KV Cache reuse. By sharing KV headers across layers, we can reduce the KV Cache size while maintaining model capacity. Specifically, the Brandon et al. (2024) method directly reuses these contents across layers, as CLA, whereas YOCO and GoldFinch in Sun et al. (2024); Goldstein et al. (2024) introduce more significant changes: they generate the all layers’ KV Cache using linear models, like RetNet (Sun et al., 2023a) and RWKV (Peng et al., 2023a). However, it’s essential to note that these methods do not optimize the memory bandwidth bottleneck of the KV Cache. Cross-layer reuse results in scattered access to the same KV Cache, and only meticulously crafted CUDA kernels can effectively reduce data exchanges between different cache levels, if such reduction is possible at all.
We can consider reuse as a form of low-rank compression, and complete low-rank compression should have even greater potential. The method proposed by DeepSeek-AI et al. (2024), called Multi-Head Latent Attention (MLA), exemplifies this. MLA achieves KV Cache decompression by utilizing an additional dimension-expanding matrix, rather than merely reusing the existing content. However, it’s essential to note that this method still does not address the memory bandwidth bottleneck, as the cache still needs to be expanded into large KV vectors.
Yen et al. (2024) proposed a new idea for the LLM architecture. CEPE, a framework that combines the pre-trained LLM with an Encoder module that serves as context compressor. In RAG scenario, the ultra-long context will greatly increase the need to store KV Cache. CEPE’s additional Encoder compresses these reference texts, and then inputs them into the Decoder through cross-attention. The compressed text has a shorter sequence length than before, so the KV Cache is saved. This method can use the pre-trained LLM, but the added decoder and cross-attention layer still require a lot of extra training.
@@ 4 Deploy-Stage Optimization
From the perspective of inference systems, KV Cache presents significant challenges in two key areas. First, during the continuous $\{K,k_{n}\}\longrightarrow K$ and $\{V,v_{n}\}\longrightarrow V$ operations, memory is repeatedly allocated and released, resulting in substantial fragmentation. Second, the inability to batch-process KV Cache exacerbates severe memory bandwidth bottlenecks, leading to computational inefficiency. Addressing these two issues has become a central concern for inference systems.
@@
Kwon et al. (2023) introduced the Paged Attention mechanism and the vLLM framework. Paged Attention draws on the page memory mechanism widely used in CPU memory, and uses an additional mapping table to map the KV Cache that used to be stored continuously to discontinuous GPU memory. During inference, in order to generate the next token, it is necessary to call KV Cache for calculation, and this process can also be efficiently completed through a set of custom CUDA kernels. The use of the Paged Attention mechanism results in almost no unused memory fragments and efficient inference.
@@
Lin et al. (2024) further developed the idea of Paged Attention. Through the DistAttention it proposed and the DistKV-LLM built on it, KV Cache was able to achieve distributed deployment on multiple servers. This significantly improved the efficiency of providing LLM services using large-scale cloud servers.
@@
Ye et al. (2024) aims to reuse KV Cache between different dialogues, achieving acceleration of the pre-fill stage and optimization of GPU memory occupancy. By establishing a dictionary tree for all historical dialogues, and finding the longest common prefix in the dictionary tree when a new dialogue request is received and reusing the corresponding KV Cache for this part, the ChunkAttention made the model avoid repeated calculation of some tokens in the pre-fill stage, speeding up the response speed of the deployment system. Moreover, Gao et al. (2024) extend this idea by introducing a hierarchical system that utilizes multiple storage devices.
Jin et al. (2024) and Lee et al. (2024) offload KV Cache to CPU with speculation. The core idea is to offload the primary portion of the KV Cache to the CPU, retaining only critical components. During inference, by speculatively reloading a small amount of KV Cache onto the GPU using the limited preserved information, we can save memory space on high-speed computational devices without compromising performance, all while avoiding extensive data exchanges.
He & Zhai (2024) introduces a novel approach: by collaboratively utilizing the CPU for attention computations, we can enhance GPU utilization efficiency, improve overall effective computational power, and accelerate inference speed.
Additionally, Qin et al. (2024) provides a comprehensive technical report from an enterprise perspective, optimizing the entire inference system.